OpenStudy (anonymous):

If log2=0.301 and log3=0.477, evaluate: log 5??

OpenStudy (yrelhan4):

log(a+b)=loga + logb can you do it now?

OpenStudy (anonymous):

not true. loga + logb = log(ab)

OpenStudy (yrelhan4):

oh damn sorry.

OpenStudy (anonymous):

euler is right

OpenStudy (yrelhan4):

haha. i messed up. apologies.

OpenStudy (anonymous):

log 5 it will use log(a/b) = loga - logv

OpenStudy (anonymous):

log b

OpenStudy (anonymous):

in my following solution i will imply that by log you mean log base 10 and not ln. correct me if im wrong

OpenStudy (anonymous):

yes you are right

OpenStudy (yrelhan4):

haha ikr eyad. :/

OpenStudy (jhannybean):

wait, is it \(\large \log _{\color{red}{2}}x=0.301\)?????

OpenStudy (anonymous):

you have to use that but

OpenStudy (jhannybean):

Then you can use what i call "flower power"

OpenStudy (yrelhan4):

log 2 to the base 10 is 0.3..

OpenStudy (jhannybean):

ohhhh thank you @yrelhan4

OpenStudy (yrelhan4):

ok what you need to do is.. log5= log(10/2) = log10 - log2.. log10=1 and you are given the balue of log 2.. solve.

OpenStudy (anonymous):

\[y = \log_{b}x ---> x = b^{y}\] so we are told that: \[0.301 = \log_{10}2 \therefore 2 = 10^{0.301}\] and \[0.477 = \log_{10}3 \therefore 3 = 10^{0.477}\] \[5 = 2+3 \therefore 5 = 10^{0.301} + 10^{0.477}\]

OpenStudy (anonymous):

yrelhan is also correct, but i think my answer is the one he expected to do since he is given both informations. both very valid methods though

OpenStudy (anonymous):

no i need to use only logs of 2 and 3

OpenStudy (yrelhan4):

Then follow what Euler said.

OpenStudy (jhannybean):



OpenStudy (anonymous):

\[\therefore\] means therefore btw if you didn tknow

OpenStudy (jhannybean):

"flower power" => 2= 10^(0.301) haha

OpenStudy (jhannybean):

exactly what @Euler271 said